sana is senior to mona by 18 years. after 6 years sana will be twice as old as mon. find mona's present age?
Answer by Maths68(1474)

Let
Present age of Sana = s
Present age of Mona = m


Sana is senior to Mona by 18 years
s=m+18............(1)

After 6 years
Sana will be = s+6
Mona will be = m+6
Then
Sana will be twice as old as Mona
(s+6)=2(m+6)
s+6=2m+12
s=2m+12-6
s=2m+6.............(2)
Put the value of s from (2) to (1)
2m+6=m+18
2m-m=18-6
m=12
Put the value of m in (1)
s=m+18
s=12+18
s=30



Present age of Sana = s = 30 years old
Present age of Mona = m = 12 years old